Help with all parts

Four 6 mu C point charges are at the corners of a
Answer #1

Use equation V= q/4πr

distance between center of square and charge = r = sqrt(5^2+5^2) = 5*sqrt2 = 7.07m   

a)

Vtotal=V1+V2+V3+V4= (1/4πr)*[q1+q2+q3+q4)

Since q1=q2=q3=q4= q=6μC

Vtotal = 4q/(4πr) = q/πr = (6*10^-6)/(3.14*7.07) = 2.7*10^-7 volt

b) when q1 is negative

q1=-6 μC

Vtotal=V1+V2+V3+V4= (1/4πr)*[q1+q2+q3-q4) = (2q)/(4πr) = 1/2*q/πr =1/2*2.7*10^-7 volt = 1.35*10^-7 volt

c) When q3 = q4= -6μC

Vtotal=V1+V2+V3+V4= (1/4πr)*[q1+q2+q3-q4) = (1/4πr)*[q+q-q-q) = 0 volt